### v2

# The expected value of sample information calculations for external validation of risk prediction models

Abstract
In designing external validation studies of clinical prediction models, contemporary sample size calculation methods are based on the frequentist inferential paradigm. One of the widely reported metrics of model performance is net benefit (NB), and the relevance of conventional inference around NB as a measure of clinical utility is doubtful. Value of Information methodology quantifies the consequences of uncertainty in terms of its impact on clinical utility of decisions. We introduce the expected value of sample information (EVSI) for validation as the expected gain in NB from conducting an external validation study of a given size. We propose algorithms for EVSI computation, and in a case study demonstrate how EVSI changes as a function of the amount of current information and future study’s sample size. Value of Information methodology provides a decision-theoretic lens to the process of planning a validation study of a risk prediction model and can complement conventional methods when designing such studies.
footnotetext: From Faculty of Medicine and Pharmaceutical Sciences (MS, TYL), and Department of Statistics (PG), The University of British Columbia, Vancouver, British Columbia, Canada; Department of Epidemiology and Biostatistics, Memorial Sloan Kettering Cancer Center, New York, New York, USA (AV); Department of Epidemiology, CAPHRI Care and Public Health Research Institute, Maastricht University, Maastricht, The Netherlands, and Department of Development and Regeneration, KU Leuven, Leuven, Belgium (LW)footnotetext: *Correspondence to Mohsen Sadatsafavi, 2405 Wesbrook Mall, Vancouver, BC, V6T1Z3, Canada; email: mohsen.sadatsafavi@ubc.ca
Introduction
Clinical prediction models that quantify the risk of medically significant events are a key aspect of individualized medicine. Before being trusted for clinical use in a target population, a model needs to undergo validation in a representative sample from that population. Model performance during such external validation is assessed in various ways, which can broadly be classified as prediction error (e.g., Brier score), discrimination (e.g., c-statistic), calibration (e.g., calibration intercept and slope), and net benefit (NB)1. Among such metrics, NB is a decision-theoretic one, as it enables direct assessment of the clinical utility of a model (i.e., if a model’s expected NB is higher than alternatives, it is expected to confer clinical utility). Due to such strong decision-theoretic underpinnings, and despite its relatively new arrival in the field of predictive analytics, NB has gained significant momentum and has become a standard component of modern external validation studies2.
When interpreting the results of a validation study, the finite size of the validation sample means that the assessment of model performance is accompanied by uncertainty. Uncertainty in conventional metrics of model performance is typically communicated using classical inferential methods (e.g., 95% confidence interval [CI] around c-statistic or calibration slope, p-values for comparing the c-statistic of competing models). However, the relevance of such frequentist approach to uncertainty around NB as a decision-theoretic metric is doubtful3–5.
Decision theory provides an alternative view to the consequences of uncertainty by relating it to the outcome of decisions. Leaving aside, for the sake of simplicity, the issue of risk aversion, the decision-maker should adopt the strategy with the highest expected utility, which, if implementation costs are comparable, is the one with the highest net benefit. While uncertainty around NB should not affect the ‘adoption decision’, it is associated with utility loss as it hinders our ability to identify the decision with the highest true NB. The extent of uncertainty should thus inform the ‘research decision’: whether the expected utility loss is large enough to necessitate collecting further evidence. This approach towards uncertainty quantification is referred to as Value of Information (VoI) analysis6–8.
VoI methods are widely accepted in health policy-making7,9. However, their relevance to clinical decision-making and risk prediction has only recently been highlighted10. In particular, the expected value of perfect information (EVPI), the expected gain in NB by completely eliminating uncertainty, has been applied to both the development and validation phases of risk prediction models10,11. If based on EVPI, the investigator concludes that uncertainty is large and procuring more validation data is warranted, the next natural question is ‘how much data should be collected?’. The expected value of sample information (EVSI) is the expected gain in NB by conducting a study of a given sample size (and design)12. In the context of model validation, EVSI can thus inform the sample size determination of future validation studies.
The aim of the present work is to define EVSI for the external validation phase of a risk prediction model and propose algorithms for its computation. The rest of this manuscript is structured as follows. After outlining the context, we define EVSI for a future study aimed at evaluating the clinical utility of a model in a target population. We propose algorithms for EVSI computation based on bootstrapping when preliminary data from the target populations are available, a beta-binomial modeling approach for binary outcomes when current information can be specified as independent Beta distributions, and a more general algorithm when the current information is expressed as an arbitrary probability distribution. A case study puts the developments in context where we provide practical suggestions on reporting and interpreting EVSI. We conclude by proposing areas of further inquiry.
Methods
Context and notations
We focus on validating a previously-developed risk prediction model in a well-defined new target population. The developments are presented for models for binary responses but are generally applicable to any context where NB can be assessed (e.g., survival outcomes13, models for treatment benefit14). The model is advertised as a function that maps patient characteristics to an estimate of event risk. We acknowledge the relevance of other sources of uncertainty (e.g., missing data, the plausibility that the sample is truly representative of the target population) but focus the developments on sampling uncertainty, and discuss some other uncertainty sources whenever relevant.
VoI analysis is a Bayesian paradigm and requires explicit specification of current information on the performance of the model in the target population. By default, we assume this information is generated from a preliminary validation exercise based on an iid sample $\mathbf{d}$ of $n$ individuals. However, such information can be generated in other ways. For example, one can extrapolate (optimism-corrected) performance of a model from its development sample, or its performance in previous validation studies15, and accommodate any added uncertainty about the compatibility of the populations (e.g. by quantitatively discounting external information16). Formal expert elicitation approaches can also be employed to construct distributions around model performance metrics17. In such instances, $\mathbf{d}$ can be an abstract entity representing, for example, the expert-elicited information.
Given our current information, we are considering the utility of procuring data of sample of size $n^{*}$ for a future validation study in the same target population. We are interested in expected utility gain from such a study. We express utility in terms of net benefit (NB), as is quantified and presented in decision curve analysis18. In a nutshell, NB is based on the specification of a risk threshold, denoted by $z$, placed on predicted risks such that individuals with predicted risk lower than this threshold are considered low-risk, and those with predicted risk higher than this threshold are classified as high-risk (and therefore are selected for treatment). For individuals with predicted risk precisely equal to $z$, the decision-maker will be ambivalent about the treatment decision. Because among these individuals, treating all means a proportion $z$ is properly treated (true positive) while the proportion $1-z$ is unnecessarily treated (false positive), such ambivalence implies that each false positive has $-z/(1-z)$ times clinical utility of a true positive. Because the true and false positive rates of the model can be estimated at this threshold, this implied exchange rate can be used to evaluate the NB of a model in net true positive or false positive units. Any given model should be compared with at least two ‘default’ strategies of treating no one and treating all. Treating no one has NB=0 by definition, and NB calculations for treating all follow the same logic as above, but by considering all individuals as positive. In practice, NB is measured over a range of plausible risk thresholds.
Let $\theta$ be the set of parameters that fully specify the target population for NB calculations. A common setup for binary outcomes is to define $\theta$ as the triplet of prevalence ($\theta_{p}$), sensitivity ($\theta_{se}$), and specificity ($\theta_{sp}$)20. We index treating no one, model-based treatment, and treating all by, respectively, subscripts 0,1,2, resulting in the following equation for NBs (for brevity, we drop the notation that indicates quantities are functions of $z$)
$NB(i,\mathbf{\theta})=\left\{\begin{array}[]{lll}0&i=0&\mbox{(treat no one)}\\ \theta_{p}\theta_{se}-(1-\theta_{p})(1-\theta_{sp})\frac{z}{1-z}&i=1&(\mbox{use model to decide})\\ \theta_{p}-(1-\theta_{p})\frac{z}{1-z}&i=2&(\mbox{treat all})\\ \end{array}\right.,$ (1)
with $\theta=\{\theta_{p},\theta_{se},\theta_{sp}\}$. For time-to-event outcomes, these quantities are time-dependent and should be defined at a time-horizon of interest13. If there are alternative strategies they can also be considered (the above indexing of $NB()$ can accommodate other strategies), but this is omitted from this work for the ease of exposition.
Our current information about $\theta$ based on the available evidence is expressed as $P(\theta|\mathbf{d})$. According to Bayes’ theorem, $P(\theta|\mathbf{d})\propto P(\theta)P(\mathbf{d}|\theta)$, indicating that this information is influenced by any prior information we might have had, as well as by what we have learned from $\mathbf{d}$.
The expected value of current information
With $P(\theta|\mathbf{d})$ quantifying their knowledge about $\theta$, a risk-neutral decision-maker should pick the strategy with the highest expected NB given the current information:
$\mathbb{E}NB_{\text{current information}}=\max_{i}{\mathbb{E}_{\mathbb{\theta}|\mathbf{d}}NB(i,\mathbb{\theta})},$ (2)
where the expectation is with respect to $P(\theta|\mathbf{d})$. Generally, evaluating this expectation might require a Monte Carlo approach (e.g., via Markov Chain Monte Carlo sampling from $P(\theta|\mathbf{d})$). However, in many instances, this expectation can be more readily obtained. For example, a Bayesian interpretation of the bootstrap allows one to interpret an estimate of $\theta$ obtained from a bootstrapped copy of $\mathbf{d}$ as a random draw from the posterior distribution of $\theta$ given $\mathbf{d}$21. However, as the number of bootstrap iterations increases, due to the law of large numbers the average values from such bootstrap samples for each component of $\theta$ as well as $NB$s converge to their point estimate in the sample. This justifies interpreting the point estimates of $NB$ from the sample, as is conducted in conventional decision curve analysis21, in a Bayesian framework as the posterior mean under a non-informative prior.
The expected value of perfect information (EVPI)
Details of the reasoning and algorithms behind calculating EVPI is presented previously11. In brief, if we could know the true values of $\theta$, we would have picked the strategy with the highest true NB. We indeed do not have access to the ground truth, but have partial information in terms of $P(\theta|\mathbf{d})$. As such, we can model the consequence of knowing the truth and take its expectation with respect to this distribution:
$\mathbb{E}NB_{\text{perfect information}}=\mathbb{E}_{\mathbf{\theta}|\mathbf{d}}\max_{i}NB(i,\mathbf{\theta}).$ (3)
The EVPI is then the difference in expected NB of having perfect information versus current information:
$EVPI=\mathbb{E}NB_{\text{perfect information}}-\mathbb{E}NB_{\text{current information}}.$ (4)
The EVPI is a scalar quantity in NB units and quantifies the expected loss in NB due to uncertainty, or alternatively, the expected gain in NB by completely resolving uncertainty (i.e., conducting a validation study of infinite size). We have proposed a generic algorithm based on bootstrapping to estimate this expectation, as well as an asymptotic approach based on central limit theorem and two-dimensional unit normal loss integral11.
The expected value of sample information (EVSI)
The reasoning behind EVSI calculation is similar to that of EVPI, with the modification that instead of knowing the truth, we will have more (but not complete) information about $\theta$, with additional information coming from a future iid sample $\mathcal{D^{*}}$ of size $n^{*}$. With this added information we will have the opportunity to revise our current decision. We will update our knowledge from $P(\theta|\mathbf{d})$ to $P(\theta|\mathbf{d},\mathcal{D^{*}})$, and will select the strategy that has the highest expected $NB$.
The NB of this approach once future data $\mathcal{D^{*}}$ are obtained is $\max_{i}\mathbb{E}_{\mathbf{\theta|{\mathcal{D^{*}},d}}}NB(i,\mathbf{\theta})$. We do not know what data we will observe in the future, but we can treat $\mathcal{D^{*}}$ as a random variable, and take the expectation of this line of reasoning with regard to its predictive distribution:
$\mathbb{E}NB_{\text{sample information}}=\mathbb{E}_{\mathcal{D^{*}}|\mathbf{d}}\max_{i}\mathbb{E}_{\mathbf{\theta}|\mathcal{D^{*}},\mathbf{d}}NB(i,\mathbb{\theta}).$ (5)
The EVSI is the difference between this ENB and ENB under current information:
$EVSI=\mathbb{E}NB_{\text{sample information}}-\mathbb{E}NB_{\text{current information}}.$ (6)
The EVSI is a non-negative scalar quantity in the same NB units as EVPI. It quantifies the expected gain in NB by procuring a further validation sample of a given size. The higher the EVSI, the higher the expected gain from the planned validation study. EVPI, quantifying the expected NB gain by completely resolving uncertainty, puts an upper limit on EVSI.
EVSI computation algorithms
The challenge for EVSI calculation is the two nested expectations. The inner expectation represents updated estimates of NBs after obtaining future data, and the outer expectation is with respect to the predictive distribution of future data. Below we detail three computaion algorithms.
A bootstrap-based algorithm
If the current evidence $\mathbf{d}$ is in terms of a random sample of size $n$ from the target population of interest, sampling from $P(\theta|\mathbf{d})$ can be performed via the Bayesian bootstrap21. The resulting algorithm for EVSI computation is similar to the previously proposed method for EVSI calculations for data-driven economic evaluations22 and can be seen as an extension of the bootstrap method for EVPI computation for external validation studies11.
The Bayesian bootstrap of $\mathbf{d}$ assigns an improper non-informative prior distribution $\text{Dirichlet}(0,0,...,0)$ across the $n$ data points. A random draw from the posterior distribution of the generating population can be materialized by drawing a random weight vector from the $\text{Dirichlet}(1,1,...,1)$ distribution. Estimating $\theta$ in this weighted sample is akin to drawing a value for $\theta$ from its posterior distribution21. Further, once the generating population is at hand, the data of the future study ($\mathcal{D^{*}}$) can be drawn by sampling from this population. This results in a two-level resampling approach for EVSI calculation explained in Table 1.
[…]
${}^{b}$ $\mathbb{E}NB_{\text{current information}}$ can also be directly evaluated in the original sample without bootstrapping. However, we recommend averaging over bootstrapped values as this approach is consistent with the way other terms are evaluated and this prevents getting occasional negative VoI metrics
Table 1: Bootstrap-based computation of EVSI
The power of this approach is in the flexibility of the bootstrap method in accommodating different types of outcomes and practical considerations in validation studies. For example, if the dataset $\mathbf{d}$ has non-trivial level of missing data, which affects the amount of information it contains, the step 1b in the above algorithm can include an imputation component, generating the population $\mathbf{F}$ without any missing values (and similarly, step 1d can introduce missingness in the data to simulate a real-world validation dataset).
The ordinary bootstrap, which assigns weights from a scaled $\text{Multinomial}(1/n,1/n,...)$ distribution to observations, is conceptually similar to the Bayesian bootstrap and has been interpreted in a Bayesian way (e.g., in the approximate Bayesian bootstrap employed in missing value imputation23). In data-driven VoI analyses it generates similar results to the Bayesian bootstrap11,22. It thus can be used in lieu of the Bayesian bootstrap in step 1a.
An algorithm based on beta-binomial conjugacy for binary outcomes
In their proposal for a parametric Bayesian NB estimation for binary outcomes, Cruz and Korthauer take advantage of the beta-binomial conjugacy and the fact that the likelihood function for $NB$ can be factorized into terms involving prevalence, sensitivity, and specificity. As such, they argue that specifying prior information in terms of independent beta distributions on these quantities will result in corresponding posterior distributions that are also independent beta20. In this case, given that $P(\theta|\mathbf{d})$ is composed of three independent beta distributions, by the same token $P(\theta|\mathbf{d},\mathcal{D^{*}})$ after a realization of $\mathcal{D^{*}}$ will also be the product of three independent beta distributions. This results in a fast Monte-Carlo approach for EVSI computation as explained in Table 2.
0. Specify the current information for NBs under current information as $\theta_{p}\sim\text{Beta}(\alpha_{p},\beta_{p})$, $\theta_{se}\sim\text{Beta}(\alpha_{se},\beta_{se})$, $\theta_{sp}\sim\text{Beta}(\alpha_{sp},\beta_{sp})$. 1. Plug in the expected values of $\theta$ from the above distributions in equation 1. Record the maximum expected NB. This is $\mathbb{E}NB_{\text{current information}}$. 2. For $j$ = 1 to $N$ (number of Monte Carlo simulations) (a) Randomly draw from the distribution of $\theta$ as specified in step 2. (b) Compute true NBs from such draws. Record the maximum NB. (c) Generate $\mathcal{D^{*}}$, defined by $\{n^{*}_{tp},n^{*}_{fn},n^{*}_{tn},n^{*}_{fp}\}$ given the draw from $\theta$ obtained in step 2a as: $n^{*}_{+}\sim\text{Binomial}(n^{*},\theta_{p})$ (number of positive cases in the future sample), $n^{*}_{tp}\sim\text{Binomial}(n^{*}_{+},\theta_{se})$, $n^{*}_{fn}=n^{*}_{+}-n^{*}_{tp}$, $n^{*}_{tn}\sim\text{Binomial}(n^{*}-n^{*}_{+},\theta_{sp})$, $n^{*}_{fp}=n^{*}-n^{*}_{+}-n^{*}_{tn}$. (d) Calculate expected NBs given generated $\mathcal{D^{*}}$ by plugging in the updated estimates of the expected value of $\theta$ as follows: ${\bar{\theta}^{*}_{p}}=(\alpha_{p}+n^{*}_{tp}+n^{*}_{fn})/(\alpha_{p}+\beta_{p}+n^{*})$, ${\bar{\theta}^{*}_{se}}=(\alpha_{se}+n^{*}_{tp})/(\alpha_{se}+\beta_{se}+n^{*}_{tp}+n^{*}_{fn})$, ${\bar{\theta}^{*}_{sp}}=(\alpha_{sp}+n^{*}_{tn})/(\alpha_{sp}+\beta_{sp}+n^{*}_{tn}+n^{*}_{fp})$. Record the maximum NB. This will be $\mathbb{E}_{\mathbf{\theta|\mathbf{d},\mathcal{D^{*}}}}NB(i,\mathbf{\theta})$. Next $j$ 3. Average NBs from step 2b. This is $\mathbb{E}NB_{\text{perfect information}}$. From this subtract ENB from step 1. This is EVPI. 4. Average NBs from step 2d. This is $\mathbb{E}NB_{\text{sample information}}$. From this subtract ENB from step 1. This is EVSI.
Table 2: Computation of EVSI based on beta-binomial conjugacy
Independent beta distributions for each component of $\theta$ can emerge, from example, if we ask an expert to express their belief about prevalence, sensitivity, and specificity, in terms of fractions. That is, if their best estimate of prevalence is 30%, and their uncertainty is akin to their opinion comping from a sample of 10 individuals, the information can be expressed as $\theta_{p}\sim\text{Beta}(3,7)$ (similar for sensitivity and specificity).
Importantly, for binary outcomes and in the absence of missing data, the Bayesian bootstrap approach explained above reduces to this beta-binomial model. This is because the information in a sample $\mathbf{d}$ can be fully specified via four quantities: the number of true positives ($n_{tp}$), false negatives ($n_{fn}$), true negatives ($n_{tn}$), and false positives ($n_{fp})$. The Bayesian bootstrapping of individual observations is akin to assigning an improper prior $\text{Dirichlet}(0,0,0,0)$ to these quantities, giving rise to the posterior distribution being of the form $\text{Dirichlet}(n_{tp},n_{fn},n_{tn},n_{fp})$. Due to the aggregate property of the Dirichlet distribution24, the $\text{Dirichlet}(0,0,0,0)$ prior of the Bayesian bootstrap is equal to assigning $\text{Beta}(0,0)$ priors to $\theta_{p}$, $\theta_{se}$, and $\theta_{sp}$. As such, the beta-binomial algorithms will be equal to the Bayesian bootstrap and can be used instead for VoI computations due to its computational efficiency.
Yet another utility of this approach is that instead of the improper prior employed in the Bayesian bootstrap, the user can specify informative priors. For example, if an independent cross-sectional study in the same target population has estimated outcome prevalence that can be expressed as $\text{Beta}(\alpha,\beta)$, then combined with observed data, our current information about prevalence can be expressed as $\theta_{p}\sim\text{Beta}(\alpha+n_{tp}+n_{fn},\beta+n_{fp}+n_{tn})$. Further, even in the absence of prior information, proper non-informative or weakly informative priors can be used to rectify some of the previously-identified problems with bootstrap-based VoI calculations at extreme situations. In small validation samples and extreme thresholds, VoI calculations might generate counter-intuitive results11. One instance is when sample estimates of $\theta_{se}$ or $\theta_{sp}$ are at their extreme (0 or 1). The improper $\text{Beta}(0,0)$ prior on sensitivity and specificity can result in improper beta posteriors with one of their two parameters being 0. Therefore, we suggest specifying proper priors (e.g., a flat prior $\text{Beta}(1,1)$) for the three quantities in $\theta$ when current information is from an observed sample.
A general algorithm for arbitrary specification of $P(\theta|\mathbf{d})$
In many practical situations our current information $P(\theta|\mathbf{d})$ may not have a simple mathematical form, but one can still draw a Monte Carlo sample of arbitrary size from its distribution (e.g., via Markov Chain Monte Carlo methods). Examples include when $P(\theta_{p})$ is constructed from a meta-analysis of prevalence studies, or $P(\theta_{se},\theta_{sp})$ from a bivariate meta-analysis of the diagnostic accuracy of the test, giving rise to a joint distribution for $P(\theta_{se},\theta_{sp})$, potentially conditional on prevalence25.
For such a general case where we have a sample from $P(\theta|\mathbf{d})$, the empirical distribution of the sample can be used as the distribution for the current information. The posterior distribution after observing future data can then be approximated by the discrete distribution that puts a mass $w^{(i)}\propto P(\mathcal{D}^{*}|\theta^{(i)})$ (e.g., a product of Binomial probabilities for binary outcomes) on each observation from this sample. This will result in the generic algorithm explained in Table 3.
0. Quantify current information given arbitrary distribution $P(\theta|\mathbf{d})$. Obtain $M$ draws $(\theta^{(1)},\theta^{(2)},...,\theta^{(M)})$ from $P(\theta|\mathbf{d})$. 1. Estimate expected NBs under current information by taking expectation of NB terms (in equation 1) with respect to the empirical distribution that puts a mass $w^{(k)}$ of $\propto 1$ on each $\theta^{(k)}$: $NB(i,\mathbf{\theta})=\left\{\begin{array}[]{ll}0&i=0\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}\theta^{(k)}_{se}-(1-\theta^{(k)}_{p})(1-\theta^{(k)}_{sp})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=1\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}-(1-\theta^{(k)}_{p})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=2\\ \end{array}\right..$ Record the maximum expected $NB$. This is $\mathbb{E}NB_{\text{current information}}$. 2. For $j$ = 1 to $N$ (number of Monte Carlo simulations) (a) Obtain a random draw from $P(\theta|\mathbf{d})$ (or loop over the draws from step 3). This is taken as the true $\theta$ in this iteration. (b) Calculate true NBs given $\theta$ obtained in the previous step. Record the maximum NB. (c) Generate $\mathcal{D^{*}}$ of size $n^{*}$, defined by $\{n^{*}_{tp},n^{*}_{fn},n^{*}_{tn},n^{*}_{fp}\}$ given the sample from $\theta$ obtained in step 2a: $n^{*}_{+}\sim\text{Binomial}(n^{*},\theta_{p})$ (number of positive cases in the future sample), $n^{*}_{tp}\sim\text{Binomial}(n^{*}_{+},\theta_{se})$, $n^{*}_{fn}=n^{*}_{+}-n^{*}_{tp}$, $n^{*}_{tn}\sim\text{Binomial}(n^{*}-n^{*}_{+},\theta_{sp})$, $n^{*}_{fp}=n^{*}-n^{*}_{+}-n^{*}_{tn}$. (d) Update $P(\theta|\mathbf{d})$ by updating the weights $w^{(k)}$ in the empirical distribution of $\theta|\mathbf{d}$ (from step 1) based on the likelihood $P(\mathcal{D^{*}}|\theta)$. For binary outcomes the weights are: $w^{(k)}\propto P_{Bin}(n^{*}_{+},n^{*},\theta_{p}^{(k)})\times P_{Bin}(n^{*}_{tp},n^{*}_{+},\theta_{se}^{(k)})\times P_{Bin}(n^{*}_{tn},n^{*}-n^{*}_{+},\theta_{sp}^{(k)})$, with $P_{Bin}(x,m,p)$ being the probability mass function of the binomial distribution at value $x$ for $m$ trials and success probability $p$. (e) Calculate updated NBs given future data by taking expectation of NB terms in step 1 using the updated weights from step 2d. Next $j$ 3. Average NBs from from step 2b. This is $\mathbb{E}NB_{\text{perfect information}}$. Subtract from this ENB from step 1. This is EVPI. 4. Average NBs from step 2e. This is $\mathbb{E}NB_{\text{sample information}}$. Subtract from this ENB from step 1. This is EVSI.
Table 3: EVSI computation for binary outcomes for general, non-conjugate distributions
Results
We used data from GUSTO-I, a clinical trial of multiple thrombolytic strategies for acute myocardial infarction (AMI), as our case study26. This dataset has been widely used for methodological research in predictive analytics27. All analyses are conducted in R28.
For the present case study, we use a similar set-up as our previous work, with a risk prediction model for 30-day mortality developed using the non-US sample of this dataset, and validated in the US sub-sample. The risk prediction model is a logit model fitted to the entire non-US subset (sample size 17796), with the resulting equation:
logit(P(Y=1))= - 2.0842 + 0.0781[age] + 0.4027[AMI location other (vs. inferior)] + 0.5773[AMI location anterior (v. inferior)] + 0.4678[previous AMI] + 0.7666[AMI severity (Killip score)] - 0.0775[min(blood pressure, 100)] + 0.0182[pulse].
Similar to our previous work, we initially assume we have access to only $n=500$ individuals from the US sub-sample as the source of current information. We randomly selected 500 individuals, without replacement, from the entire 23034 observations in the US sub-sample of GUSTO-I. Later, we will use more observations from the US sub-sample to study how EVSI behaves with the amount of current information. The prevalence of the outcome in the development sample (entire non-US sub-sample) was 0.0723; in the current validation sample it was 0.0860, while in the entire US sub-sample it was 0.0679. We consider the range of relevant thresholds to be 1%-10%.
We start by drawing the decision curve at this range of thresholds, as shown in Figure 1. Panel (a) shows the decision curve, and panel (b) shows the difference between the NB of the model and the best alternative default strategy ($NB_{1}-\max(NB_{0},NB_{2})$) along with its 95% confidence interval (CI, based on the percentile methods using 10,000 bootstraps). The model had higher expected NB than the alternative strategies at the entire range of thresholds of interest.
(a) Decision curve
(b) Incremental NB of the model v. best alternative strategy
Figure 1: Decision curve (a) and incremental net benefit of the model v. the best alternative strategy (b). Dashed lines are 95% confidence intervals
Next, we draw the EVPI curve. EVPI and EVSI computations are based on 10^6 Monte Carlo simulations using the aggregate beta-binomial model, with a flat $\text{Beta}(1,1)$ prior on each component of $\theta$. The EVPI at 0.01 threshold was 0.00039. At the 0.02 threshold it was 0.00125. The EVPI was 0 at 0.05 and 0.10 thresholds (so EVSI for any $n^{*}$ will also be 0). We therefore focus on VoI calculations at 0.01 and 0.02 thresholds.
The EVPI curve is presented in Figure 2. EVPI quantifies the expected NB loss per decision (every time the model might be used). The overall NB loss due to uncertainty is therefore affected by the expected number of times the medical decision will be made29. In scaling VoI quantities to the population, we apply a similar approach as in our previous work11: a decision-making body in charge of making a national recommendation can consider the model to be applicable to all AMI cases in the US (the target population of the validation sample). There are around 800,000 AMIs in the US every year30, and a guideline development panel in charge of recommending risk stratification for AMI management can consider all those instances relevant. The scaled VoI values are provided in true positive units on the second Y-axis of the EVPI figure. At the 0.02 threshold, a future validation study can have a maximum benefit equal to gaining 1001 net true positive (individuals who will die within 30 days and are correctly identified as high risk) per year, or avoiding 49069 extra false positive cases per year (individuals who will not die within 30 days.
Figure 2: The expected value of perfect information (EVPI) of the validation sample
The EVSI curves are shown in Figure 3. Generally, it is expected that EVSI will increase with a larger future study, and will asymptote to EVPI, a pattern that is obvious for both thresholds. The ‘diminishing return’ pattern is also obvious for both thresholds: the expected NB gain steeply rises at small samples and plateaus as $n^{*}$ grows. Similar to EVPI, population scaling is applied to EVSIs and is presented as the second Y-axis of the EVSI curve. At 0.02 threshold, a future study of size $n^{*}=500$ has a per-decision EVSI value of 0.00100, corresponding to population value of 804 in true positive cases gained, or 39381 in false positive cases averted.
(a) $z=0.01$
(b) $z=0.02$
Figure 3: EVSI for the case study
Figure 4 provides the results of brief simulation studies that demonstrate how EVSI changes as a function of the amount of current information, represented by the sample size of $\mathbf{d}$, the current validation data from which $P(\theta|\mathbf{d})$ is constructed. Starting from a value of $500$, we doubled the sample size of the current validation sample, specified the distributions for $\theta_{p}$, $\theta_{se}$, and $\theta_{sp}$ based on this sample, and performed EVSI calculations as above, until reaching $n=8,000$. Results are provided for both 0.01 and 0.02 thresholds and are the average of 100 independent simulations, in each a new sample (with replacement) from the US subset of GUSTO-I was obtained as the current validation data (100 simulations were considered adequate based on inspecting the standard error of the Monte Carlo simulations).
As expected, the higher the amount of current information (the larger the $n$), the lower the gain in NB by conducting future studies. When the current validation sample is small, the initial gain in NB by conducting future research can be very large, as evidenced by the steep slope of the EVSI curve. This is not the case when the current information is relatively robust (e.g., based on $n$=8,000, which includes on average 544 events). The comparison between the two thresholds is also informative, as it demonstrates that the amount of gain from future sample is highly dependent on the threshold, even in relative terms. For example, at $z=0.01$, the EVSI curve when current information is based on $n$=1,000 (dotted blue line) is only slightly lower than the EVSI curve when $n$=500 (dotted black line). At $z=0.02$, EVSI when current information is based on $n$=1,000 (solid blue line) is significantly lower than when the current information is based on $n$=500 (solid black line), across the entire range of future sample sizes.
Figure 4: The expected value of sample information (EVSI) as a function of the sample sample size of the future study ($\mathcal{D^{*}}$) for various levels of current information, represented by the sample size of the current study ($n$) from which $P(\theta)$ is constructed
Discussion
Evaluating the performance of a model in a finite sample is fraught with uncertainty. In this work, we applied the VoI framework to investigate the decision-theoretic consequences of such uncertainty. We defined validation EVSI as the expected gain in NB by obtaining an external validation sample of a given size from the target population of interest. In a case study we showed the feasibility of EVSI calculations, and studied how EVSI is affected by the amount of current information and the sample size of the future study. We suggested scaling the EVSI to the population to quantify the overall gain in clinical utility in true (or false) positive units from conducting an external validation study of a given sample size.
How can EVSI analysis inform study design in predictive analytics? Conducting a validation study is an investment in resources that will generate further information on the performance of a clinical prediction model in a target population. The decision whether to undertake a validation study should ultimately hinge on whether the information gained from such a study is worth the required investments. The EVSI, when scaled to the population, determines the expected return on investment in NB unit. Ultimately, such return should be contrasted against the efforts and resources required for such a study. In decision-analytic (health policy) modeling, EVSI is typically in net monetary units, and when scaled to population, can be compared with the budget of a planned data collection activity. The optimal sample size will be one that maximizes the difference between population EVSI and study costs29. The NB for risk prediction models, on the other hand, is in net true or false positive units, and as such cannot directly be compared with the budget of a validation study. One can always embark on full decision analysis to translate all outcomes to net monetary units, but this will likely require sophisticated decision modeling and context-specific assumptions on long-term outcomes, a process that might take significant amount of time and require further data collection (e.g., to obtain utility weights for outcomes). To us, a main reason for the vast popularity of decision curve analysis is that it provides an assumption-free, reproducible method for NB calculation based on the very same data that are used for studying model calibration and discrimination. We think VoI analysis during model development and validation should generally keep the same spirit. A full decision analysis should be relegated to after an impact analysis has measured the resource-use implications of implementing the model. This, however, means the decision rules for determining the optimal sample size of development and validation studies based on the VoI framework would be different than those used in health policy analysis. This paper deliberately stayed away from proposing such rules, and instead focused on defining concepts and proposing computation methods for EVSI. Proposing such decision rules is detached from EVSI calculations and deserves its own airing.
There are multiple areas of further inquiry. The EVSI framework should also be applied to the development phase of prediction models. This can guide the investigator on whether further development, or moving to validation, should be prioritized. We mainly focused on NB loss due to sampling uncertainty. However, there are several sources of uncertainty, such as whether our existing information on the performance of the model is directly applicable to the target population, or if predictors and outcome are measured with the same quality between the study and usual practice. The comparative statistical and computational performance of the EVSI computations algorithms, and the adequacy of a given number of simulations for each algorithm, should be evaluated in dedicated studies. We also do not claim the algorithms we proposed are the only ones that can be used for EVSI computations. Other algorithms, such as those based on central limit theorem31,32, can prove useful. VoI analysis in decision modeling has received significant boost in computational speed in recent years due to the arrival of algorithms based on non-parametric regression modeling33,34. This approach can facilitate VoI analysis in risk prediction as well. The EVSI defined in this work is for a single, homogeneous target population, and does not consider heterogeneous settings. Consider the hierarchical Bayesian tri-variate model by Wynants et al for NB estimation in multi-center validation studies19 and imagine a new center is interested in validating this model in their local setting. The posterior predictive distribution from the hierarchical model for a new center (potentially combined with any local information on the performance of the model) can act as ‘current information’, enabling EVSI calculations. This will be valid and informative insofar as the expected NB gain in that center is concerned. However, in a hierarchical Bayesian framework, information learned from one cluster propagates to others, and thus conducting a validation study in a new center results in change in the distribution of NB in other centers. Such NB gain from learning across clusters in hierarchical settings is not reflected in the EVSI computed for this particular setting. VoI metrics and corresponding computation algorithms for multi-center studies should be developed separately. Further, during external validation, often a secondary aim is to update the model if its performance turns out to be sub-optimal35. Such model revision can take different levels of complexity (from intercept correction to re-estimating regression coefficients36) and one might be interested in the expected yield of a given sample size for such model revision. This will get connected to the VoI concepts for model development. As stated earlier, how VoI metrics for prediction models should inform objective functions for determining the optimal sample size should be debated by the community in the hope of generating consensus and best practice standards.
There is an ongoing debate on the appropriateness of conventional, frequentist metrics of uncertainty when reporting NB of a clinical prediction model3–5. The same concerns can logically be extended to frequentist method for sample size and power calculations around NB37–39. VoI methodology provides a rigorous, utilitarian response to such controversies. The toolbox of VoI methods for clinical prediction models is growing, and perhaps it is time to formalize the role of VoI in uncertainty quantification and design of empirical studies in predictive analytics.

### v3

@@
 Introduction
-Clinical prediction models that quantify the risk of medically significant events are a key aspect of individualized medicine.
-Before being trusted for clinical use in a target population, a model needs to undergo validation in a representative sample from that population.
-Model performance during such external validation is assessed in various ways, which can broadly be classified as prediction error (e.g., Brier score), discrimination (e.g., c-statistic), calibration (e.g., calibration intercept and slope), and net benefit (NB)1.
-Among such metrics, NB is a decision-theoretic one, as it enables direct assessment of the clinical utility of a model (i.e., if a model’s expected NB is higher than alternatives, it is expected to confer clinical utility).
-Due to such strong decision-theoretic underpinnings, and despite its relatively new arrival in the field of predictive analytics, NB has gained significant momentum and has become a standard component of modern external validation studies2.
+Risk prediction models that quantify the probability of clinical events are a key aspect of individualized medicine.
+Once developed, these models are often used in diverse populations.
+Before being trusted for clinical use in a new population, a prediction model needs to undergo validation in a representative sample from that population1.
+For example, the original Framingham risk scores for cardiovascular disease risk are developed using data from the Framingham County in the US, but are externally validated for use in many different populations2.
+During a validation study, the performance of the prediction model is assessed in terms of average prediction error, discrimination (e.g., c-statistic), calibration (e.g., calibration intercept and slope), and net benefit (NB)3.
+Among such metrics, NB is a decision-theoretic one, as it enables direct assessment of the clinical utility of a model (i.e., if a model’s expected NB is higher than that of alternatives, it is expected to confer clinical utility).
+Due to such decision-theoretic underpinnings, and despite its relatively new arrival in the field of predictive analytics, NB has gained significant momentum and has become a standard component of modern external validation studies4.
 When interpreting the results of a validation study, the finite size of the validation sample means that the assessment of model performance is accompanied by uncertainty.
-Uncertainty in conventional metrics of model performance is typically communicated using classical inferential methods (e.g., 95% confidence interval [CI] around c-statistic or calibration slope, p-values for comparing the c-statistic of competing models).
-However, the relevance of such frequentist approach to uncertainty around NB as a decision-theoretic metric is doubtful3–5.
+Uncertainty in conventional metrics of model performance is communicated using classical inferential methods (e.g., 95% confidence interval [CI] around c-statistic or calibration slope).
+Similarly, sample size considerations when planning a validation study are based on inferential statistics5,6.
+However, the relevance of this approach to uncertainty around NB as a decision-theoretic metric is doubtful, as arbitrary significance levels or pre-specified confidence bands do not necessarily translate to better decisions7–9.
 Decision theory provides an alternative view to the consequences of uncertainty by relating it to the outcome of decisions.
-Leaving aside, for the sake of simplicity, the issue of risk aversion, the decision-maker should adopt the strategy with the highest expected utility, which, if implementation costs are comparable, is the one with the highest net benefit.
+A decision-maker should adopt the strategy with the highest expected utility, which, if implementation costs are comparable, is the one with the highest expected NB.
 While uncertainty around NB should not affect the ‘adoption decision’, it is associated with utility loss as it hinders our ability to identify the decision with the highest true NB.
 The extent of uncertainty should thus inform the ‘research decision’: whether the expected utility loss is large enough to necessitate collecting further evidence.
-This approach towards uncertainty quantification is referred to as Value of Information (VoI) analysis6–8.
-VoI methods are widely accepted in health policy-making7,9.
-However, their relevance to clinical decision-making and risk prediction has only recently been highlighted10.
-In particular, the expected value of perfect information (EVPI), the expected gain in NB by completely eliminating uncertainty, has been applied to both the development and validation phases of risk prediction models10,11.
-If based on EVPI, the investigator concludes that uncertainty is large and procuring more validation data is warranted, the next natural question is ‘how much data should be collected?’.
-The expected value of sample information (EVSI) is the expected gain in NB by conducting a study of a given sample size (and design)12.
-In the context of model validation, EVSI can thus inform the sample size determination of future validation studies.
-The aim of the present work is to define EVSI for the external validation phase of a risk prediction model and propose algorithms for its computation.
+This approach towards uncertainty quantification is referred to as Value of Information (VoI) analysis10–12.
+VoI methods are widely accepted in health policy-making11,13.
+However, their relevance to clinical decision-making and risk prediction has only recently been highlighted14.
+In particular, the expected value of perfect information (EVPI), the expected gain in NB by completely eliminating uncertainty, has been applied to both the development and validation phases of risk prediction models14,15.
+However, to the best of our knowledge, the expected value of sample information (EVSI), the expected gain in NB by conducting a study of a given sample size, has not been defined and applied to clinical prediction models.
+The aim of the present work is to define EVSI for the external validation studies of risk prediction models and propose algorithms for its computation.
 The rest of this manuscript is structured as follows.
-After outlining the context, we define EVSI for a future study aimed at evaluating the clinical utility of a model in a target population.
-We propose algorithms for EVSI computation based on bootstrapping when preliminary data from the target populations are available, a beta-binomial modeling approach for binary outcomes when current information can be specified as independent Beta distributions, and a more general algorithm when the current information is expressed as an arbitrary probability distribution.
-A case study puts the developments in context where we provide practical suggestions on reporting and interpreting EVSI.
+After outlining the context, we define EVSI for a future validation study aimed at evaluating the clinical utility of a model in a new population.
+We propose different algorithms for EVSI computations, elaborate on their use case, and compare their performance in simulation studies.
+A case study puts the developments in context.
 We conclude by proposing areas of further inquiry.
 Methods
-Context and notations
-We focus on validating a previously-developed risk prediction model in a well-defined new target population.
-The developments are presented for models for binary responses but are generally applicable to any context where NB can be assessed (e.g., survival outcomes13, models for treatment benefit14).
-The model is advertised as a function that maps patient characteristics to an estimate of event risk.
-We acknowledge the relevance of other sources of uncertainty (e.g., missing data, the plausibility that the sample is truly representative of the target population) but focus the developments on sampling uncertainty, and discuss some other uncertainty sources whenever relevant.
-VoI analysis is a Bayesian paradigm and requires explicit specification of current information on the performance of the model in the target population.
-By default, we assume this information is generated from a preliminary validation exercise based on an iid sample $\mathbf{d}$ of $n$ individuals.
-However, such information can be generated in other ways.
-For example, one can extrapolate (optimism-corrected) performance of a model from its development sample, or its performance in previous validation studies15, and accommodate any added uncertainty about the compatibility of the populations (e.g.
-by quantitatively discounting external information16).
-Formal expert elicitation approaches can also be employed to construct distributions around model performance metrics17.
-In such instances, $\mathbf{d}$ can be an abstract entity representing, for example, the expert-elicited information.
-Given our current information, we are considering the utility of procuring data of sample of size $n^{*}$ for a future validation study in the same target population.
-We are interested in expected utility gain from such a study.
-We express utility in terms of net benefit (NB), as is quantified and presented in decision curve analysis18.
-In a nutshell, NB is based on the specification of a risk threshold, denoted by $z$, placed on predicted risks such that individuals with predicted risk lower than this threshold are considered low-risk, and those with predicted risk higher than this threshold are classified as high-risk (and therefore are selected for treatment).
-For individuals with predicted risk precisely equal to $z$, the decision-maker will be ambivalent about the treatment decision.
-Because among these individuals, treating all means a proportion $z$ is properly treated (true positive) while the proportion $1-z$ is unnecessarily treated (false positive), such ambivalence implies that each false positive has $-z/(1-z)$ times clinical utility of a true positive.
-Because the true and false positive rates of the model can be estimated at this threshold, this implied exchange rate can be used to evaluate the NB of a model in net true positive or false positive units.
+Context
+We focus on validating a previously-developed risk prediction model in a new target population.
+The developments are presented for prediction models for binary responses but are generally applicable to any context where NB can be assessed (e.g., survival outcomes16, models for treatment benefit17).
+A risk prediction model is advertised as a function that maps patient characteristics to an estimate of event risk.
+We have some ‘current’ information about the performance of this model in the target population (for example, from expert opinion, or from a pilot study), but are uncertain if using the model in this population is net beneficial.
+We are planning to obtain a random sample $\mathcal{D}$ containing $\mathit{N}$ observations from this population to reduce such uncertainty.
+This sample is expected to improve our knowledge about model performance, which in turn will increase our chance of making the correct decision whether to use the model in this population.
+As such, procuring such a sample might be associated with NB gain, and we are interested in evaluating the expected gain in NB as a function of $\mathit{N}$.
+We acknowledge the relevance of other sources of uncertainty (e.g., missing data, the plausibility that the sample is truly representative of the target population) but focus on sampling uncertainty – that is, our uncertainty in true NBs because of the finite size of the validation sample.
+Net benefit calculations for risk prediction models
+Details of NB calculations for risk prediction models are explained in its original paper and in many tutorials4,18,19.
+In brief, to turn a continuous predicted risk to a binary classification to inform a treatment decision, a decision-maker needs to specify a risk threshold $z$ on predicted risks such that individuals classified as high-risk are selected for the treatment (e.g., a 7.5%-10% threshold on Framingham risk score is often used to initiate treatment with statins to reduce cardiovascular disease risk20).
+Compared to treating no one (with a default utility of 0), the utility of this classification can be expressed as $p_{tp}-\omega p_{fp}$, where $p_{tp}$ and $p_{fp}$ are the probabilities of true positive and false positives, respectively, at this threshold, and $\omega\geq 0$ represents the utility tradeoff (exchange rate) between a true positive and a false positive classification.
+Vickers and Elkin showed that the value of $\omega$ can be deduced from the risk threshold itself18: if a decision-maker is willing to treat those with risk $>z$ and not treat those with risk $<z$, it means among individuals with risk precisely equal to $z$, they will be ambivalent about the treatment decision.
+This ambivalence indicates that in a population whose average outcome risk is $z$, the expected utility of treating everyone and treating no one is the same to the decision-maker.
+Because in this population, treating everyone corresponds to $p_{tp}=z$ and $p_{fp}=1-z$, we have $\omega=z/(1-z)$.
+For example, a 10% threshold on cardiovascular risk for initiating statin therapy means the benefit of treating one individual who will experience a cardiovascular event is considered equal to the harm of treating nine individuals who will not experience such an event without treatment.
+This exchange rate enables calculating NB in either net true positive or net false positive units for a given threshold.
+In practice, the NB is calculated across a plausible range of thresholds.
+It is often more intuitive to express $p_{tp}$ and $p_{fp}$ in terms of outcome prevalence and the model’s true sensitivity and specificity in the target population21.
+Let $\theta$ be the triplet of outcome prevalence ($\theta_{p}$, a quantity that is independent of model performance), sensitivity ($\theta_{se}$), and specificity ($\theta_{sp}$).
+We have $p_{tp}=\theta_{p}\theta_{se}$ and $p_{fp}=(1-\theta_{p})(1-\theta_{sp})$ (for brevity, we drop the notation that indicates sensitivity and specificity are functions of $z$).
 Any given model should be compared with at least two ‘default’ strategies of treating no one and treating all.
 Treating no one has NB=0 by definition, and NB calculations for treating all follow the same logic as above, but by considering all individuals as positive.
-In practice, NB is measured over a range of plausible risk thresholds.
-Let $\theta$ be the set of parameters that fully specify the target population for NB calculations.
-A common setup for binary outcomes is to define $\theta$ as the triplet of prevalence ($\theta_{p}$), sensitivity ($\theta_{se}$), and specificity ($\theta_{sp}$)20.
-We index treating no one, model-based treatment, and treating all by, respectively, subscripts 0,1,2, resulting in the following equation for NBs (for brevity, we drop the notation that indicates quantities are functions of $z$)
+Thus we have three competing strategies: do not treat anyone, use the model to treat those with predicted risk $\geq z$, and treat all.
+We index them by 0,1,2, respectively.
+This results in the following equation for NBs:
 $NB(i,\mathbf{\theta})=\left\{\begin{array}[]{lll}0&i=0&\mbox{(treat no one)}\\ \theta_{p}\theta_{se}-(1-\theta_{p})(1-\theta_{sp})\frac{z}{1-z}&i=1&(\mbox{use model to decide})\\ \theta_{p}-(1-\theta_{p})\frac{z}{1-z}&i=2&(\mbox{treat all})\\ \end{array}\right.,$ (1)
-with $\theta=\{\theta_{p},\theta_{se},\theta_{sp}\}$.
-For time-to-event outcomes, these quantities are time-dependent and should be defined at a time-horizon of interest13.
-If there are alternative strategies they can also be considered (the above indexing of $NB()$ can accommodate other strategies), but this is omitted from this work for the ease of exposition.
+For time-to-event outcomes, $\theta$s and thus NBs are time-dependent and should be defined at a time-horizon of interest16.
+If there are alternative strategies, they can also be considered (the above indexing of NB() can accommodate other strategies), but this is omitted here for the ease of exposition.
+Value of Information analysis
+VoI analysis is a Bayesian approach and requires explicit specification of current information on the performance of the model in the target population.
+By default, we assume this information is generated from a preliminary validation exercise based on a random sample $\mathbf{d}$ of size $n$.
+However, such information can be generated in other ways.
+For example, one can extrapolate the performance of a model from its development sample, or its performance in previous validation studies22, and accommodate any added uncertainty about the compatibility of the populations by quantitatively discounting external information23.
+Formal expert elicitation approaches can also be employed to construct distributions around model performance metrics24.
+In such instances, $\mathbf{d}$ can be an abstract entity representing, for example, the expert-elicited information.
+The expected value of current information
 Our current information about $\theta$ based on the available evidence is expressed as $P(\theta|\mathbf{d})$.
-According to Bayes’ theorem, $P(\theta|\mathbf{d})\propto P(\theta)P(\mathbf{d}|\theta)$, indicating that this information is influenced by any prior information we might have had, as well as by what we have learned from $\mathbf{d}$.
-The expected value of current information
-With $P(\theta|\mathbf{d})$ quantifying their knowledge about $\theta$, a risk-neutral decision-maker should pick the strategy with the highest expected NB given the current information:
-$\mathbb{E}NB_{\text{current information}}=\max_{i}{\mathbb{E}_{\mathbb{\theta}|\mathbf{d}}NB(i,\mathbb{\theta})},$ (2)
+According to Bayes’ theorem, $P(\theta|\mathbf{d})\propto P(\theta)P(\mathbf{d}|\theta)$, indicating that this information is influenced by any prior information we might have had, as well as by $P(\mathbf{d}|\theta)$, i.e., the likelihood function, the support the observed data provide for a given value of $\theta$.
+With such current information, the decision-maker should pick the strategy with the highest expected NB given the current information:
+${{E}}NB_{\text{current}}=\max_{{}_{i\in\{0,1,2\}}}{\mathbb{E}_{\mathbb{\theta}|\mathbf{d}}NB(i,\mathbb{\theta})},$ (2)
 where the expectation is with respect to $P(\theta|\mathbf{d})$.
-Generally, evaluating this expectation might require a Monte Carlo approach (e.g., via Markov Chain Monte Carlo sampling from $P(\theta|\mathbf{d})$).
-However, in many instances, this expectation can be more readily obtained.
-For example, a Bayesian interpretation of the bootstrap allows one to interpret an estimate of $\theta$ obtained from a bootstrapped copy of $\mathbf{d}$ as a random draw from the posterior distribution of $\theta$ given $\mathbf{d}$21.
-However, as the number of bootstrap iterations increases, due to the law of large numbers the average values from such bootstrap samples for each component of $\theta$ as well as $NB$s converge to their point estimate in the sample.
-This justifies interpreting the point estimates of $NB$ from the sample, as is conducted in conventional decision curve analysis21, in a Bayesian framework as the posterior mean under a non-informative prior.
 The expected value of perfect information (EVPI)
-Details of the reasoning and algorithms behind calculating EVPI is presented previously11.
-In brief, if we could know the true values of $\theta$, we would have picked the strategy with the highest true NB.
+Details of the reasoning and algorithms behind calculating EVPI are presented previously15.
+In brief, if we could know the true values of $\theta$, we would pick the strategy with the highest true NB.
 We indeed do not have access to the ground truth, but have partial information in terms of $P(\theta|\mathbf{d})$.
-As such, we can model the consequence of knowing the truth and take its expectation with respect to this distribution:
-$\mathbb{E}NB_{\text{perfect information}}=\mathbb{E}_{\mathbf{\theta}|\mathbf{d}}\max_{i}NB(i,\mathbf{\theta}).$ (3)
+As such, we can model the consequence of knowing the truth (having perfect information) and take its expectation with respect to this distribution:
+${{E}}NB_{\text{truth}}=\mathbb{E}_{\mathbf{\theta}|\mathbf{d}}\left[\max_{{}_{i\in\{0,1,2\}}}NB(i,\mathbf{\theta})\right].$ (3)
 The EVPI is then the difference in expected NB of having perfect information versus current information:
-$EVPI=\mathbb{E}NB_{\text{perfect information}}-\mathbb{E}NB_{\text{current information}}.$ (4)
-The EVPI is a scalar quantity in NB units and quantifies the expected loss in NB due to uncertainty, or alternatively, the expected gain in NB by completely resolving uncertainty (i.e., conducting a validation study of infinite size).
-We have proposed a generic algorithm based on bootstrapping to estimate this expectation, as well as an asymptotic approach based on central limit theorem and two-dimensional unit normal loss integral11.
+$EVPI={{E}}NB_{\text{truth}}-{{E}}NB_{\text{current}}.$ (4)
+The EVPI is a scalar quantity and quantifies the expected loss in NB due to uncertainty, or alternatively, the expected gain in NB by completely resolving uncertainty (i.e., conducting a validation study of infinite size).
+We have proposed a generic algorithm based on bootstrapping to estimate this expectation, as well as an asymptotic approach based on central limit theorem and two-dimensional unit normal loss integral15.
 The expected value of sample information (EVSI)
-The reasoning behind EVSI calculation is similar to that of EVPI, with the modification that instead of knowing the truth, we will have more (but not complete) information about $\theta$, with additional information coming from a future iid sample $\mathcal{D^{*}}$ of size $n^{*}$.
+The reasoning behind EVSI calculation is similar to that of EVPI, with the modification that instead of knowing the truth, we will have more (but not perfect) information about $\theta$, with additional information coming from obtaining a future sample $\mathcal{D}$.
 With this added information we will have the opportunity to revise our current decision.
-We will update our knowledge from $P(\theta|\mathbf{d})$ to $P(\theta|\mathbf{d},\mathcal{D^{*}})$, and will select the strategy that has the highest expected $NB$.
-The NB of this approach once future data $\mathcal{D^{*}}$ are obtained is $\max_{i}\mathbb{E}_{\mathbf{\theta|{\mathcal{D^{*}},d}}}NB(i,\mathbf{\theta})$.
-We do not know what data we will observe in the future, but we can treat $\mathcal{D^{*}}$ as a random variable, and take the expectation of this line of reasoning with regard to its predictive distribution:
-$\mathbb{E}NB_{\text{sample information}}=\mathbb{E}_{\mathcal{D^{*}}|\mathbf{d}}\max_{i}\mathbb{E}_{\mathbf{\theta}|\mathcal{D^{*}},\mathbf{d}}NB(i,\mathbb{\theta}).$ (5)
+We will update our knowledge from $P(\theta|\mathbf{d})$ to $P(\theta|\mathbf{d},\mathcal{D})$, and will select the strategy that has the highest expected NB.
+The NB of this approach once future data are obtained is $\max_{i}\mathbb{E}_{\mathbf{\theta|{\mathcal{D},d}}}NB(i,\mathbf{\theta})$.
+We do not know what data we will observe in the future, but we can treat $\mathcal{D}$ as a random variable, and take the expectation of this line of reasoning with regard to its predictive distribution (the distribution for future observations given current data and the model):
+${{E}}NB_{\text{sample}}=\mathbb{E}_{\mathcal{D}|\mathbf{d}}\left[\max_{{}_{i\in\{0,1,2\}}}\mathbb{E}_{\mathbf{\theta}|\mathbf{d},\mathcal{D}}NB(i,\mathbb{\theta})\right].$ (5)
 The EVSI is the difference between this ENB and ENB under current information:
-$EVSI=\mathbb{E}NB_{\text{sample information}}-\mathbb{E}NB_{\text{current information}}.$ (6)
+$EVSI={{E}}NB_{\text{sample}}-{{E}}NB_{\text{current}}.$ (6)
 The EVSI is a non-negative scalar quantity in the same NB units as EVPI.
-It quantifies the expected gain in NB by procuring a further validation sample of a given size.
+It quantifies the expected gain in NB by procuring a validation sample of a given size.
 The higher the EVSI, the higher the expected gain from the planned validation study.
@@
 EVSI computation algorithms
-The challenge for EVSI calculation is the two nested expectations.
+The challenge for EVSI calculation is the two nested expectations in ${{E}}NB_{\text{sample}}$.
 The inner expectation represents updated estimates of NBs after obtaining future data, and the outer expectation is with respect to the predictive distribution of future data.
-Below we detail three computaion algorithms.
+Below we detail three computation algorithms.
 A bootstrap-based algorithm
-If the current evidence $\mathbf{d}$ is in terms of a random sample of size $n$ from the target population of interest, sampling from $P(\theta|\mathbf{d})$ can be performed via the Bayesian bootstrap21.
-The resulting algorithm for EVSI computation is similar to the previously proposed method for EVSI calculations for data-driven economic evaluations22 and can be seen as an extension of the bootstrap method for EVPI computation for external validation studies11.
-The Bayesian bootstrap of $\mathbf{d}$ assigns an improper non-informative prior distribution $\text{Dirichlet}(0,0,...,0)$ across the $n$ data points.
-A random draw from the posterior distribution of the generating population can be materialized by drawing a random weight vector from the $\text{Dirichlet}(1,1,...,1)$ distribution.
-Estimating $\theta$ in this weighted sample is akin to drawing a value for $\theta$ from its posterior distribution21.
-Further, once the generating population is at hand, the data of the future study ($\mathcal{D^{*}}$) can be drawn by sampling from this population.
-This results in a two-level resampling approach for EVSI calculation explained in Table 1.
-[…]
-${}^{b}$ $\mathbb{E}NB_{\text{current information}}$ can also be directly evaluated in the original sample without bootstrapping.
-However, we recommend averaging over bootstrapped values as this approach is consistent with the way other terms are evaluated and this prevents getting occasional negative VoI metrics
+If the current evidence $\mathbf{d}$ is in terms of a random sample of size $n$ from the target population (e.g., from a pilot study), EVSI computation can be performed via a two-level resampling algorithm.
+This approach has been previously applied to EVSI computation for data-driven economic evaluations25 and can be seen as a direct extension of the bootstrap method for EVPI computation for external validation studies15.
+The logic behind this approach is provided in detail previously25.
+In a nutshell, let $\mathcal{D^{*}}$ be the target population from which $\mathbf{d}$ and $\mathcal{D}$ are sampled.
+We take $\mathcal{D^{*}}$ as a random entity and apply the Bayes’s rule: $P(\mathcal{D^{*}}|\mathbf{d})\propto P(\mathcal{D^{*}})P(\mathbf{d}|\mathcal{D^{*}})$.
+Rubin showed that by using the improper prior $P(\mathcal{D^{*}})\sim\text{Dirichlet}(0,0,…,0)$ placed on all possible observations in $\mathcal{D^{*}}$, the posterior distribution $P(\mathcal{D^{*}}|\mathbf{d})$ will be a discrete distribution that puts random weights Dirichlet(1,1,…,1) on each observation in $\mathbf{d}$.
+Since any parameter of interest $\theta$ is a function of $\mathcal{D^{*}}$, posterior draws of weights from the $\text{Dirichlet}(1,1,…,1)$ distribution produce draws from the posterior distribution of $P(\theta|\mathbf{d})$26.
+Further, once the generating population is at hand, the data of the future study ($\mathcal{D}$) can be drawn by sampling from this population.
+The resulting two-level resampling algorithm is explained in Table 1.
+- Prerequisite: Obtain $\mathbf{d}$, a pilot sample of size $n$ consisting of predicted risks and observed responses from the target population.
+1. For $j$ = 1 to $M$ (number of Monte Carlo simulations)
+(a) Draw $\mathcal{D^{*}}$, a Bayesian bootstrap of the data $\mathbf{d}$, as a random draw from the posterior distribution of the population that has generated the sample.
+This is done by assigning random weights $W^{*}\sim\text{Dirichlet}(1,1,...,1)$ to each observation in $\mathbf{d}$.
+(b) Estimate $\theta^{*}$, the true values of prevalence, sensitivity, and specificity in this iteration, from $\mathcal{D^{*}}$.
+(c) Calculate $NB^{*}_{0}$, $NB^{*}_{1}$, and $NB^{*}_{2}$, by plugging $\theta^{*}$ into Equation 1.
+These are true NBs in this iteration.
+Record the maximum of true NBs: $NB^{*}_{\text{truth}}=\max\{NB^{*}_{0},NB^{*}_{1},NB^{*}_{2}\}$.
+(d) Draw $\mathcal{D}$, a random realization of the data of the future study, through sampling $\mathit{N}$ observations with replacement from $\mathcal{D^{*}}$.
+(e) Create $\mathcal{D^{+}}=\mathbf{d}+\mathcal{D}$, the pooled data of existing and future samples.
+(f) Calculate $\theta^{+}$ from $\mathcal{D^{+}}$, the updated estimates of prevalence, sensitivity, and specificity after obtaining the future sample.
+(g) Calculate $NB^{+}_{0}$, $NB^{+}_{1}$, and $NB^{+}_{2}$, updated estimates of expected NBs after obtaining the future sample, by plugging $\theta^{+}$ into Equation 1.
+Determine the strategy that has the highest expected future NB: $l={\mbox{argmax}_{i\in\{0,1,2\}}(NB^{+}_{i})}$.
+Record the true NB of this strategy: $NB^{*}_{\text{sample}}=NB^{*}_{l}$.
+- Next $j$
+2. Take the average of $NB^{*}$s from step 1c and pick the maximum value. This is ${{E}}NB_{\text{current}}$§.
+3. Average $NB^{*}_{\text{truth}}$s from step 1c. This is ${{E}}NB_{\text{truth}}$. From this subtract ${{E}}NB_{\text{current}}$. This is EVPI.
+4. Average $NB^{*}_{\text{sample}}$s from step 1g. This is ${{E}}NB_{\text{sample}}$. From this subtract ${{E}}NB_{\text{current}}$. This is EVSI.
+Ordinary bootstrap can also be used - see text.
+One way of generating such wights is by generating $n$ exponential(1) random variables and scaling them to add up to 1.
+§ Because the NB terms are linear on the elements of $\theta$, ${{E}}NB_{\text{current}}$ can also be directly evaluated in the original sample without bootstrapping, by plugging in the point estimates of $\theta$ into Equation 1.
+However, we recommend averaging over bootstrapped values as the positive correlation between this term and the other term in EVPI and EVSI equations improves precision and prevents getting occasional negative VoI estimates.
 Table 1: Bootstrap-based computation of EVSI
 The power of this approach is in the flexibility of the bootstrap method in accommodating different types of outcomes and practical considerations in validation studies.
-For example, if the dataset $\mathbf{d}$ has non-trivial level of missing data, which affects the amount of information it contains, the step 1b in the above algorithm can include an imputation component, generating the population $\mathbf{F}$ without any missing values (and similarly, step 1d can introduce missingness in the data to simulate a real-world validation dataset).
-The ordinary bootstrap, which assigns weights from a scaled $\text{Multinomial}(1/n,1/n,...)$ distribution to observations, is conceptually similar to the Bayesian bootstrap and has been interpreted in a Bayesian way (e.g., in the approximate Bayesian bootstrap employed in missing value imputation23).
-In data-driven VoI analyses it generates similar results to the Bayesian bootstrap11,22.
+For example, if $\mathbf{d}$ has non-trivial level of missing data, which affects the amount of information it contains, Step 1c in Table 1 can include an imputation component, generating the population $\mathcal{D^{*}}$ without any missing values (and similarly, Step 1d can introduce missingness in the future sample to simulate a real-world validation dataset).
+The ordinary bootstrap, which assigns weights from a scaled $\text{Multinomial}(1,1,...,1)$ distribution to observations in $\mathbf{d}$, is conceptually similar to the Bayesian bootstrap and has been interpreted in a Bayesian way (e.g., in the approximate Bayesian bootstrap employed in missing value imputation27).
+In data-driven VoI analyses it generates similar results to the Bayesian bootstrap25.
 It thus can be used in lieu of the Bayesian bootstrap in step 1a.
-An algorithm based on beta-binomial conjugacy for binary outcomes
-In their proposal for a parametric Bayesian NB estimation for binary outcomes, Cruz and Korthauer take advantage of the beta-binomial conjugacy and the fact that the likelihood function for $NB$ can be factorized into terms involving prevalence, sensitivity, and specificity.
-As such, they argue that specifying prior information in terms of independent beta distributions on these quantities will result in corresponding posterior distributions that are also independent beta20.
-In this case, given that $P(\theta|\mathbf{d})$ is composed of three independent beta distributions, by the same token $P(\theta|\mathbf{d},\mathcal{D^{*}})$ after a realization of $\mathcal{D^{*}}$ will also be the product of three independent beta distributions.
-This results in a fast Monte-Carlo approach for EVSI computation as explained in Table 2.
-0.
-Specify the current information for NBs under current information as $\theta_{p}\sim\text{Beta}(\alpha_{p},\beta_{p})$, $\theta_{se}\sim\text{Beta}(\alpha_{se},\beta_{se})$, $\theta_{sp}\sim\text{Beta}(\alpha_{sp},\beta_{sp})$.
-1.
-Plug in the expected values of $\theta$ from the above distributions in equation 1.
-Record the maximum expected NB.
-This is $\mathbb{E}NB_{\text{current information}}$.
-2.
-For $j$ = 1 to $N$ (number of Monte Carlo simulations) (a) Randomly draw from the distribution of $\theta$ as specified in step 2.
-(b) Compute true NBs from such draws.
-Record the maximum NB.
-(c) Generate $\mathcal{D^{*}}$, defined by $\{n^{*}_{tp},n^{*}_{fn},n^{*}_{tn},n^{*}_{fp}\}$ given the draw from $\theta$ obtained in step 2a as: $n^{*}_{+}\sim\text{Binomial}(n^{*},\theta_{p})$ (number of positive cases in the future sample), $n^{*}_{tp}\sim\text{Binomial}(n^{*}_{+},\theta_{se})$, $n^{*}_{fn}=n^{*}_{+}-n^{*}_{tp}$, $n^{*}_{tn}\sim\text{Binomial}(n^{*}-n^{*}_{+},\theta_{sp})$, $n^{*}_{fp}=n^{*}-n^{*}_{+}-n^{*}_{tn}$.
-(d) Calculate expected NBs given generated $\mathcal{D^{*}}$ by plugging in the updated estimates of the expected value of $\theta$ as follows: ${\bar{\theta}^{*}_{p}}=(\alpha_{p}+n^{*}_{tp}+n^{*}_{fn})/(\alpha_{p}+\beta_{p}+n^{*})$, ${\bar{\theta}^{*}_{se}}=(\alpha_{se}+n^{*}_{tp})/(\alpha_{se}+\beta_{se}+n^{*}_{tp}+n^{*}_{fn})$, ${\bar{\theta}^{*}_{sp}}=(\alpha_{sp}+n^{*}_{tn})/(\alpha_{sp}+\beta_{sp}+n^{*}_{tn}+n^{*}_{fp})$.
-Record the maximum NB.
-This will be $\mathbb{E}_{\mathbf{\theta|\mathbf{d},\mathcal{D^{*}}}}NB(i,\mathbf{\theta})$.
-Next $j$ 3.
-Average NBs from step 2b.
-This is $\mathbb{E}NB_{\text{perfect information}}$.
-From this subtract ENB from step 1.
-This is EVPI.
-4.
-Average NBs from step 2d.
-This is $\mathbb{E}NB_{\text{sample information}}$.
-From this subtract ENB from step 1.
-This is EVSI.
-Table 2: Computation of EVSI based on beta-binomial conjugacy
-Independent beta distributions for each component of $\theta$ can emerge, from example, if we ask an expert to express their belief about prevalence, sensitivity, and specificity, in terms of fractions.
+A fast algorithm based on beta-binomial modeling for binary outcomes
+In their proposal for a parametric Bayesian NB estimation for binary outcomes, Cruz and Korthauer note that the likelihood function for $NB$ can be factorized into terms involving prevalence, sensitivity, and specificity28.
+As such, they take advantage of the beta-binomial conjugacy and note that specifying prior information in terms of independent beta distributions on these quantities will result in corresponding posterior distributions that are also independent beta.
+In this case, given that $P(\theta|\mathbf{d})$ is composed of three independent beta distributions, by the same token $P(\theta|\mathbf{d},\mathcal{D})$ after a realization of $\mathcal{D}$ will also be composed of three independent beta distributions.
+This results in a fast Monte Carlo approach for EVSI computation as explained in Table 2.
+- Prerequisite: Specify current information $P(\theta|\mathbf{d})$ as
+$\theta_{p}\sim\text{Beta}(\alpha_{p},\beta_{p})$, $\theta_{se}\sim\text{Beta}(\alpha_{se},\beta_{se})$, $\theta_{sp}\sim\text{Beta}(\alpha_{sp},\beta_{sp})$.
+1. For $j$ = 1 to $M$ (number of Monte Carlo simulations)
+(a) Obtain $\theta^{*}$, by random drawing from $P(\theta|\mathbf{d})$. These are true $\theta$s in this iteration.
+(b) Calculate $NB^{*}_{0}$, $NB^{*}_{1}$, and $NB^{*}_{2}$, by plugging $\theta^{*}$ into Equation 1.
+These are true NBs in this iteration.
+Record the maximum true NB: $NB^{*}_{\text{truth}}=\max\{NB^{*}_{0},NB^{*}_{1},NB^{*}_{2}\}$.
+(c) Generate $\mathcal{D}$, the sample of future study, defined by $\{\mathit{N}_{tp},\mathit{N}_{fn},\mathit{N}_{tn},\mathit{N}_{fp}\}$ given $\theta^{*}$ obtained in Step 1a as:
+$\mathit{N}_{+}\sim\text{Binomial}(N,\theta_{p}^{*})$ (number of positive cases in the future sample),
+$\mathit{N}_{tp}\sim\text{Binomial}(\mathit{N}_{+},\theta_{se}^{*})$,
+$\mathit{N}_{fn}=\mathit{N}_{+}-\mathit{N}_{tp}$,
+$\mathit{N}_{tn}\sim\text{Binomial}(\mathit{N}-\mathit{N}_{+},\theta_{sp}^{*})$,
+$\mathit{N}_{fp}=\mathit{N}-\mathit{N}_{+}-\mathit{N}_{tn}$.
+(d) Calculate $\theta^{+}$, the updated estimates of prevalence, sensitivity, and specificity after obtaining the future sample. :
+${\theta^{+}_{p}}=(\alpha_{p}+\mathit{N}_{tp}+\mathit{N}_{fn})/(\alpha_{p}+\beta_{p}+\mathit{N})$,
+${\theta^{+}_{se}}=(\alpha_{se}+\mathit{N}_{tp})/(\alpha_{se}+\beta_{se}+\mathit{N}_{tp}+\mathit{N}_{fn})$,
+${\theta^{+}_{sp}}=(\alpha_{sp}+\mathit{N}_{tn})/(\alpha_{sp}+\beta_{sp}+\mathit{N}_{tn}+\mathit{N}_{fp})$.
+(e) Calculate $NB^{+}_{0}$, $NB^{+}_{1}$, and $NB^{+}_{2}$, updated estimates of NBs after obtaining the future sample, by plugging $\theta^{+}$ into Equation 1.
+Determine the strategy that has the highest expected future NB: $l={\mbox{argmax}_{i\in\{0,1,2\}}(NB^{+}_{i})}$.
+Record the true NB of this strategy: $NB^{*}_{\text{sample}}=NB^{*}_{l}$.
+- Next $j$
+2. Take the average of $NB^{*}$s from step 1b and pick the maximum value. This is ${{E}}NB_{\text{current}}$.
+3. Average $NB^{*}_{\text{truth}}$s from step 1b. This is ${{E}}NB_{\text{truth}}$. From this subtract ${{E}}NB_{\text{current}}$. This is EVPI.
+4. Average $NB^{*}_{\text{sample}}$s from step 1e. This is ${{E}}NB_{\text{sample}}$. From this subtract ${{E}}NB_{\text{current}}$. This is EVSI.
+Table 2: Computation of EVSI based on beta-binomial modeling
+Independent beta distributions for each component of $\theta$ can emerge, for example, if we ask an expert to express their belief about prevalence, sensitivity, and specificity, in terms of fractions.
 That is, if their best estimate of prevalence is 30%, and their uncertainty is akin to their opinion comping from a sample of 10 individuals, the information can be expressed as $\theta_{p}\sim\text{Beta}(3,7)$ (similar for sensitivity and specificity).
-Importantly, for binary outcomes and in the absence of missing data, the Bayesian bootstrap approach explained above reduces to this beta-binomial model.
-This is because the information in a sample $\mathbf{d}$ can be fully specified via four quantities: the number of true positives ($n_{tp}$), false negatives ($n_{fn}$), true negatives ($n_{tn}$), and false positives ($n_{fp})$.
-The Bayesian bootstrapping of individual observations is akin to assigning an improper prior $\text{Dirichlet}(0,0,0,0)$ to these quantities, giving rise to the posterior distribution being of the form $\text{Dirichlet}(n_{tp},n_{fn},n_{tn},n_{fp})$.
-Due to the aggregate property of the Dirichlet distribution24, the $\text{Dirichlet}(0,0,0,0)$ prior of the Bayesian bootstrap is equal to assigning $\text{Beta}(0,0)$ priors to $\theta_{p}$, $\theta_{se}$, and $\theta_{sp}$.
-As such, the beta-binomial algorithms will be equal to the Bayesian bootstrap and can be used instead for VoI computations due to its computational efficiency.
+Importantly, for binary outcomes and in the absence of missing data, the Bayesian bootstrap approach explained previously reduces to this beta-binomial model.
+This is because the information in a sample $\mathbf{d}$ for estimating NBs can be fully specified via four quantities: the number of true positives ($n_{tp}$), false negatives ($n_{fn}$), true negatives ($n_{tn}$), and false positives ($n_{fp})$.
+Due to the aggregate property of the Dirichlet distribution29, the $\text{Dirichlet}(0,0,...,0)$ prior used in the Bayesian bootstrap is equal to assigning $\text{Beta}(0,0)$ priors to $\theta_{p}$, $\theta_{se}$, and $\theta_{sp}$, resulting in the posterior distribution for the population that can be specified as $(p_{tp},p_{fn},p_{tn},p_{fp})\sim\text{Dirichlet}(n_{tp},n_{fn},n_{tn},n_{fp})$.
+As such, the beta-binomial algorithm will be equal to the Bayesian bootstrap and can be used instead for VoI computations due to its computational efficiency.
 Yet another utility of this approach is that instead of the improper prior employed in the Bayesian bootstrap, the user can specify informative priors.
-For example, if an independent cross-sectional study in the same target population has estimated outcome prevalence that can be expressed as $\text{Beta}(\alpha,\beta)$, then combined with observed data, our current information about prevalence can be expressed as $\theta_{p}\sim\text{Beta}(\alpha+n_{tp}+n_{fn},\beta+n_{fp}+n_{tn})$.
-Further, even in the absence of prior information, proper non-informative or weakly informative priors can be used to rectify some of the previously-identified problems with bootstrap-based VoI calculations at extreme situations.
-In small validation samples and extreme thresholds, VoI calculations might generate counter-intuitive results11.
+Imagine in the current sample $\mathbf{d}$ of size $n$, there are $m$ individuals who have experienced the outcome.
+If an independent cross-sectional study in the same target population has estimated outcome prevalence that can be expressed as $\text{Beta}(\alpha,\beta)$, our current information about prevalence can be expressed as $\theta_{p}\sim\text{Beta}(\alpha+m,\beta+n-m)$.
+Even in the absence of external information, weakly informative priors can be employed to rectify some of the previously-identified problems with bootstrap-based VoI calculations at extreme situations.
+In small validation samples and extreme thresholds, VoI calculations might generate counter-intuitive results15.
 One instance is when sample estimates of $\theta_{se}$ or $\theta_{sp}$ are at their extreme (0 or 1).
-The improper $\text{Beta}(0,0)$ prior on sensitivity and specificity can result in improper beta posteriors with one of their two parameters being 0.
-Therefore, we suggest specifying proper priors (e.g., a flat prior $\text{Beta}(1,1)$) for the three quantities in $\theta$ when current information is from an observed sample.
+The improper $\text{Beta}(0,0)$ prior on these quantities results in improper beta posteriors with one of their two parameters being 0.
+This can be resolved if one uses a flat $\text{Beta}(1,1)$ prior for these quantities.
 A general algorithm for arbitrary specification of $P(\theta|\mathbf{d})$
-In many practical situations our current information $P(\theta|\mathbf{d})$ may not have a simple mathematical form, but one can still draw a Monte Carlo sample of arbitrary size from its distribution (e.g., via Markov Chain Monte Carlo methods).
-Examples include when $P(\theta_{p})$ is constructed from a meta-analysis of prevalence studies, or $P(\theta_{se},\theta_{sp})$ from a bivariate meta-analysis of the diagnostic accuracy of the test, giving rise to a joint distribution for $P(\theta_{se},\theta_{sp})$, potentially conditional on prevalence25.
+In many practical situations our current information $P(\theta|\mathbf{d})$ may not have a simple mathematical form, but one can still draw a Monte Carlo sample of arbitrary size from its distribution.
+Examples include when $P(\theta_{p})$ is constructed from a meta-analysis of prevalence studies, or $P(\theta_{se},\theta_{sp})$ from a bivariate meta-analysis of the diagnostic accuracy of the test, giving rise to a joint distribution for $P(\theta_{se},\theta_{sp})$30.
+Markov-Chain Monte-Carlo methods can be used to obtain samples from the joint distribution of $\theta$, even though the mathematical form of the joint density might be intractable (an example is the case studies in Wynants et al21).
 For such a general case where we have a sample from $P(\theta|\mathbf{d})$, the empirical distribution of the sample can be used as the distribution for the current information.
-The posterior distribution after observing future data can then be approximated by the discrete distribution that puts a mass $w^{(i)}\propto P(\mathcal{D}^{*}|\theta^{(i)})$ (e.g., a product of Binomial probabilities for binary outcomes) on each observation from this sample.
-This will result in the generic algorithm explained in Table 3.
-0.
-Quantify current information given arbitrary distribution $P(\theta|\mathbf{d})$.
-Obtain $M$ draws $(\theta^{(1)},\theta^{(2)},...,\theta^{(M)})$ from $P(\theta|\mathbf{d})$.
-1.
-Estimate expected NBs under current information by taking expectation of NB terms (in equation 1) with respect to the empirical distribution that puts a mass $w^{(k)}$ of $\propto 1$ on each $\theta^{(k)}$: $NB(i,\mathbf{\theta})=\left\{\begin{array}[]{ll}0&i=0\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}\theta^{(k)}_{se}-(1-\theta^{(k)}_{p})(1-\theta^{(k)}_{sp})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=1\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}-(1-\theta^{(k)}_{p})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=2\\ \end{array}\right..$ Record the maximum expected $NB$.
-This is $\mathbb{E}NB_{\text{current information}}$.
-2.
-For $j$ = 1 to $N$ (number of Monte Carlo simulations) (a) Obtain a random draw from $P(\theta|\mathbf{d})$ (or loop over the draws from step 3).
-This is taken as the true $\theta$ in this iteration.
-(b) Calculate true NBs given $\theta$ obtained in the previous step.
-Record the maximum NB.
-(c) Generate $\mathcal{D^{*}}$ of size $n^{*}$, defined by $\{n^{*}_{tp},n^{*}_{fn},n^{*}_{tn},n^{*}_{fp}\}$ given the sample from $\theta$ obtained in step 2a: $n^{*}_{+}\sim\text{Binomial}(n^{*},\theta_{p})$ (number of positive cases in the future sample), $n^{*}_{tp}\sim\text{Binomial}(n^{*}_{+},\theta_{se})$, $n^{*}_{fn}=n^{*}_{+}-n^{*}_{tp}$, $n^{*}_{tn}\sim\text{Binomial}(n^{*}-n^{*}_{+},\theta_{sp})$, $n^{*}_{fp}=n^{*}-n^{*}_{+}-n^{*}_{tn}$.
-(d) Update $P(\theta|\mathbf{d})$ by updating the weights $w^{(k)}$ in the empirical distribution of $\theta|\mathbf{d}$ (from step 1) based on the likelihood $P(\mathcal{D^{*}}|\theta)$.
-For binary outcomes the weights are: $w^{(k)}\propto P_{Bin}(n^{*}_{+},n^{*},\theta_{p}^{(k)})\times P_{Bin}(n^{*}_{tp},n^{*}_{+},\theta_{se}^{(k)})\times P_{Bin}(n^{*}_{tn},n^{*}-n^{*}_{+},\theta_{sp}^{(k)})$, with $P_{Bin}(x,m,p)$ being the probability mass function of the binomial distribution at value $x$ for $m$ trials and success probability $p$.
-(e) Calculate updated NBs given future data by taking expectation of NB terms in step 1 using the updated weights from step 2d.
-Next $j$ 3.
-Average NBs from from step 2b.
-This is $\mathbb{E}NB_{\text{perfect information}}$.
-Subtract from this ENB from step 1.
-This is EVPI.
-4.
-Average NBs from step 2e.
-This is $\mathbb{E}NB_{\text{sample information}}$.
-Subtract from this ENB from step 1.
-This is EVSI.
+We note that $P(\theta|\mathbf{d},\mathcal{D})\propto P(\theta)P(\mathbf{d},\mathcal{D}|\theta)=P(\theta)P(\mathbf{d}|\theta)P(\mathcal{D}|\theta)\propto P(\theta|\mathbf{d})P(\mathcal{D}|\theta)$.
+As such, the posterior distribution of $\theta$ after observing the future sample can be approximated by the discrete distribution that puts a mass $w^{(i)}\propto P(\mathcal{D}|\theta^{(i)})$ on the $i$th observation of the sample from $P(\theta|\mathbf{d})$.
+This will result in the general algorithm explained in Table 3.
+- Prerequisite: Specify current information as arbitrary distribution $P(\theta|\mathbf{d})$. Obtain $M$ draws $(\theta^{(1)},\theta^{(2)},...,\theta^{(M)})$ from $P(\theta|\mathbf{d})$.
+1. For $j$ = 1 to $M$ (number of Monte Carlo simulations)
+(a) Let $\theta^{*}=\theta^{(i)}$. This is true $\theta$ in this $\dagger$.
+(b) Calculate $NB^{*}_{0}$, $NB^{*}_{1}$, and $NB^{*}_{2}$, by plugging $\theta^{*}$ into Equation 1.
+These are true NBs in this iteration.
+Record the maximum true NB: $NB^{*}_{\text{truth}}=\max\{NB^{*}_{0},NB^{*}_{1},NB^{*}_{2}\}$.
+(c) Generate $\mathcal{D}$, the sample of future study, defined by $\{\mathit{N}_{tp},\mathit{N}_{fn},\mathit{N}_{tn},\mathit{N}_{fp}\}$ as follows:
+$\mathit{N}_{+}\sim\text{Binomial}(\mathit{N},\theta^{*}_{p})$ (number of positive cases in the future sample),
+$\mathit{N}_{tp}\sim\text{Binomial}(\mathit{N}_{+},\theta^{*}_{se})$,
+$\mathit{N}_{fn}=\mathit{N}_{+}-\mathit{N}_{tp}$,
+$\mathit{N}_{tn}\sim\text{Binomial}(\mathit{N}-\mathit{N}_{+},\theta^{*}_{sp})$,
+$\mathit{N}_{fp}=\mathit{N}-\mathit{N}_{+}-\mathit{N}_{tn}$.
+(d) Update the distribution of $\theta$ given the future sample. This is done by updating the weights $w^{(k)}$ based on the likelihood $P(\mathcal{D}|\theta)$. For binary outcomes the weights are:
+$w^{(k)}\propto P_{Bin}(\mathit{N}_{tp}+\mathit{N}_{fn},\mathit{N},\theta_{p}^{(k)})\times P_{Bin}(\mathit{N}_{tp},\mathit{N}_{tp}+\mathit{N}_{fn},\theta_{se}^{(k)})\times P_{Bin}(\mathit{N}_{tn},\mathit{N}_{tn}+\mathit{N}_{fp},\theta_{sp}^{(k)})$,
+with $P_{Bin}(x,m,p)$ being the probability mass function of the binomial distribution at value $x$ for $m$ trials and success probability $p$.
+(e) Calculate $NB^{+}_{0}$, $NB^{+}_{1}$, and $NB^{+}_{2}$, updated estimates of expected NBs after obtaining the future sample, by using the weights generated in the previous step as follows:
+$NB^{+}_{i}=\left\{\begin{array}[]{ll}0&i=0\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}\theta^{(k)}_{se}-(1-\theta^{(k)}_{p})(1-\theta^{(k)}_{sp})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=1\\ \sum_{k=1}^{M}w^{(k)}\{\theta^{(k)}_{p}-(1-\theta^{(k)}_{p})\frac{z}{1-z}\}/\sum_{k=1}^{M}w^{(k)}&i=2\\ \end{array}\right..$
+Determine the strategy that has the highest expected NB in this pooled sample: $l={\mbox{argmax}_{i\in\{0,1,2\}}(NB^{+}_{i})}$.
+Record the true NB of this strategy: $NB^{*}_{\text{sample}}=NB^{*}_{l}$.
+- Next $j$
+2. Average $NB^{*}_{\text{truth}}$s from from step 1b. This is ${{E}}NB_{\text{truth}}$. Subtract from this ${{E}}NB_{\text{current}}$. This is EVPI.
+3. Average $NB^{*}_{\text{sample}}$s from step 1e. This is ${{E}}NB_{\text{sample}}$. Subtract from this ${{E}}NB_{\text{current}}$. This is EVSI.
+$\dagger$ If the size of the sample is large it might not be computationally feasible to loop over all observations.
+In which case one can pick one of $\theta^{(i)}s$ at random in this step, but still uses the entire sample in step 1e.
 Table 3: EVSI computation for binary outcomes for general, non-conjugate distributions
+Case Study
+We used data from GUSTO-I, a clinical trial of multiple thrombolytic strategies for acute myocardial infarction (AMI), as our case study31.
+This dataset has been widely used for methodological research in predictive analytics32–34.
+All analyses are conducted in R35, with a fast implementation of the general EVSI algorithm in C++ (please refer to the evsiexval R package for details).
+For the present case study, we use a similar setup as our previous work, with a risk prediction model for 30-day mortality developed using the non-US sample of this dataset, and validated in the US sub-sample.
+The risk prediction model is a logistic regression model fitted to the entire non-US subset (sample size 17,796).
+We initially assume we have access to only $n=500$ individuals from the US sub-sample as the source of current information.
+We randomly selected 500 individuals, without replacement, from the entire 23,034 observations in the US sub-sample of GUSTO-I.
+Later, we will use more observations from the US sub-sample to study how EVSI behaves with the amount of current information.
+We consider the range of relevant thresholds to be 1%-10%, with 1%, 2%, 5%, and 10% being of primary interest.
+VoI computations are based on 10^6 Monte Carlo simulations using the aggregate beta-binomial model, with a flat $\text{Beta}(1,1)$ prior on each component of $\theta$.
+Using this setup, we performed two simulation studies to explore the face validity of the proposed algorithms.
+In the first set, we aimed at comparing the numerical stability and computational time of the three algorithms.
+In the second set, we explored how EVSI changes as a function of the amount of current information, represented by $n$, the size of the current validation data from which $P(\theta|\mathbf{d})$ is constructed.
 Results
-We used data from GUSTO-I, a clinical trial of multiple thrombolytic strategies for acute myocardial infarction (AMI), as our case study26.
-This dataset has been widely used for methodological research in predictive analytics27.
-All analyses are conducted in R28.
-For the present case study, we use a similar set-up as our previous work, with a risk prediction model for 30-day mortality developed using the non-US sample of this dataset, and validated in the US sub-sample.
-The risk prediction model is a logit model fitted to the entire non-US subset (sample size 17796), with the resulting equation:
+The prevalence of the outcome in the development sample (entire non-US sub-sample) was 0.0723; in the current validation sample it was 0.0860, while in the entire US sub-sample it was 0.0679.
+The risk prediction model was of the form:
 logit(P(Y=1))= - 2.0842 + 0.0781[age] + 0.4027[AMI location other (vs.
@@
 inferior)] + 0.4678[previous AMI] + 0.7666[AMI severity (Killip score)] - 0.0775[min(blood pressure, 100)] + 0.0182[pulse].
-Similar to our previous work, we initially assume we have access to only $n=500$ individuals from the US sub-sample as the source of current information.
-We randomly selected 500 individuals, without replacement, from the entire 23034 observations in the US sub-sample of GUSTO-I.
-Later, we will use more observations from the US sub-sample to study how EVSI behaves with the amount of current information.
-The prevalence of the outcome in the development sample (entire non-US sub-sample) was 0.0723; in the current validation sample it was 0.0860, while in the entire US sub-sample it was 0.0679.
-We consider the range of relevant thresholds to be 1%-10%.
-We start by drawing the decision curve at this range of thresholds, as shown in Figure 1.
-Panel (a) shows the decision curve, and panel (b) shows the difference between the NB of the model and the best alternative default strategy ($NB_{1}-\max(NB_{0},NB_{2})$) along with its 95% confidence interval (CI, based on the percentile methods using 10,000 bootstraps).
+The decision curve depicting the NB of the model alongside alternative strategies is presented in Figure 1.
+Panel (a) shows the NB, and panel (b) shows the difference between the NB of the model and the best alternative default strategy ($NB_{1}-\max(NB_{0},NB_{2})$) along with its 95% CI (based on the percentile methods using 10,000 bootstraps).
 The model had higher expected NB than the alternative strategies at the entire range of thresholds of interest.
-(a) Decision curve
-(b) Incremental NB of the model v. best alternative strategy
-Figure 1: Decision curve (a) and incremental net benefit of the model v. the best alternative strategy (b). Dashed lines are 95% confidence intervals
-Next, we draw the EVPI curve.
-EVPI and EVSI computations are based on 10^6 Monte Carlo simulations using the aggregate beta-binomial model, with a flat $\text{Beta}(1,1)$ prior on each component of $\theta$.
-The EVPI at 0.01 threshold was 0.00039.
+(a) Decision curve. Solid blue: use the model; dashed gray: treat all; dotted gray: treat none
+(b) Incremental NB of the model versus best alternative strategy
+Figure 1: Decision curve (a) and incremental net benefit of the model versus the best alternative strategy (b). Dashed lines are 95% confidence intervals
+Among 500 observations in the current sample, 43 experienced the outcome.
+Given the prior $\text{Beta}(1,1)$, this will result in $\theta_{p}\sim\text{Beta}(44,458)$.
+The distribution of sensitivity and specificity depends on the threshold.
+Taking $z=0.02$ as an example, given that 41 of the 43 individuals who experienced the outcome had predicted risk $\geq z$, and 147 of the 457 who did not experience the outcome had predicted risk $<z$, with a $\text{Beta}(1,1)$ prior, we have $\theta_{se}\sim\text{Beta}(42,3)$ and $\theta_{se}\sim\text{Beta}(148,311)$.
+The EVPI curve is presented in Figure 2.
+The EVPI at 0.01 threshold was 0.00037.
 At the 0.02 threshold it was 0.00125.
-The EVPI was 0 at 0.05 and 0.10 thresholds (so EVSI for any $n^{*}$ will also be 0).
+The EVPI was 0 at 0.05 and 0.10 thresholds (so EVSI for any $\mathit{N}$ will also be 0).
 We therefore focus on VoI calculations at 0.01 and 0.02 thresholds.
-The EVPI curve is presented in Figure 2.
 EVPI quantifies the expected NB loss per decision (every time the model might be used).
-The overall NB loss due to uncertainty is therefore affected by the expected number of times the medical decision will be made29.
-In scaling VoI quantities to the population, we apply a similar approach as in our previous work11: a decision-making body in charge of making a national recommendation can consider the model to be applicable to all AMI cases in the US (the target population of the validation sample).
-There are around 800,000 AMIs in the US every year30, and a guideline development panel in charge of recommending risk stratification for AMI management can consider all those instances relevant.
+The overall NB loss due to uncertainty is therefore affected by the expected number of times the medical decision will be made36.
+In scaling VoI quantities to the population, we apply a similar approach as in our previous work15: There are around 800,000 AMIs in the US every year37, and a national guideline development panel in charge of recommending risk stratification for AMI management can consider all those instances relevant.
 The scaled VoI values are provided in true positive units on the second Y-axis of the EVPI figure.
-At the 0.02 threshold, a future validation study can have a maximum benefit equal to gaining 1001 net true positive (individuals who will die within 30 days and are correctly identified as high risk) per year, or avoiding 49069 extra false positive cases per year (individuals who will not die within 30 days.
+At the 0.02 threshold, a future validation study can have a maximum benefit equal to gaining 1,004 net true positive (individuals who will die within 30 days and are correctly identified as high risk) per year, or avoiding 49,188 extra false positive cases per year (individuals who will not die within 30 days but are identified as high risk).
 Figure 2: The expected value of perfect information (EVPI) of the validation sample
@@
 Generally, it is expected that EVSI will increase with a larger future study, and will asymptote to EVPI, a pattern that is obvious for both thresholds.
-The ‘diminishing return’ pattern is also obvious for both thresholds: the expected NB gain steeply rises at small samples and plateaus as $n^{*}$ grows.
+The ‘diminishing return’ pattern is also clear: the expected NB gain steeply rises at small samples and plateaus as $\mathit{N}$ grows, leaving little room for gaining NB beyond $\mathit{N}$=4,000.
 Similar to EVPI, population scaling is applied to EVSIs and is presented as the second Y-axis of the EVSI curve.
-At 0.02 threshold, a future study of size $n^{*}=500$ has a per-decision EVSI value of 0.00100, corresponding to population value of 804 in true positive cases gained, or 39381 in false positive cases averted.
-(a) $z=0.01$
-(b) $z=0.02$
-Figure 3: EVSI for the case study
-Figure 4 provides the results of brief simulation studies that demonstrate how EVSI changes as a function of the amount of current information, represented by the sample size of $\mathbf{d}$, the current validation data from which $P(\theta|\mathbf{d})$ is constructed.
-Starting from a value of $500$, we doubled the sample size of the current validation sample, specified the distributions for $\theta_{p}$, $\theta_{se}$, and $\theta_{sp}$ based on this sample, and performed EVSI calculations as above, until reaching $n=8,000$.
-Results are provided for both 0.01 and 0.02 thresholds and are the average of 100 independent simulations, in each a new sample (with replacement) from the US subset of GUSTO-I was obtained as the current validation data (100 simulations were considered adequate based on inspecting the standard error of the Monte Carlo simulations).
-As expected, the higher the amount of current information (the larger the $n$), the lower the gain in NB by conducting future studies.
-When the current validation sample is small, the initial gain in NB by conducting future research can be very large, as evidenced by the steep slope of the EVSI curve.
-This is not the case when the current information is relatively robust (e.g., based on $n$=8,000, which includes on average 544 events).
-The comparison between the two thresholds is also informative, as it demonstrates that the amount of gain from future sample is highly dependent on the threshold, even in relative terms.
-For example, at $z=0.01$, the EVSI curve when current information is based on $n$=1,000 (dotted blue line) is only slightly lower than the EVSI curve when $n$=500 (dotted black line).
-At $z=0.02$, EVSI when current information is based on $n$=1,000 (solid blue line) is significantly lower than when the current information is based on $n$=500 (solid black line), across the entire range of future sample sizes.
-Figure 4: The expected value of sample information (EVSI) as a function of the sample sample size of the future study ($\mathcal{D^{*}}$) for various levels of current information, represented by the sample size of the current study ($n$) from which $P(\theta)$ is constructed
+At 0.02 threshold, a future study of size $\mathit{N}=1,000$ has a per-decision EVSI value of 0.00101, corresponding to population value of 806 in true positive cases gained, or 39,500 in false positive cases averted.
+Figure 3: The expected value of sample information (EVSI) for the case study.
+Red: EVSI; blue (horizontal): the expected value of perfect information (EVPI).
+Dashed lines: $z$ (risk threshold)=0.01; solid lines: $z$=0.02
 Discussion
@@
 We suggested scaling the EVSI to the population to quantify the overall gain in clinical utility in true (or false) positive units from conducting an external validation study of a given sample size.
+We proposed three algorithms for EVSI computations (with implementation in R as part of the evsiexval package: https://github.com/resplab/evsiexval).
+The bootstrap-based algorithm is applicable when previous individual-level data (e.g., from a pilot validation study) are at hand.
+The main advantage of this algorithm is in its flexibility, for example in mimicking expected patterns of missingness in data.
+Among the three algorithms, this algorithm is the one that can most readily be extended to survival outcomes (similar to the extension of the bootstrap method for inference around NB16).
+However, this algorithm is applicable only when individual-level data are available and can be slow.
+The beta-binomial algorithm is applicable when current evidence on model performance can be expressed as independent beta distributions for prevalence, sensitivity, and specificity, and is the fastest of the three algorithms.
+We showed that for binary outcomes and in the absence of censoring this algorithm is equal to the Bayesian bootstrap (with the added flexibility that prior information can be incorporated).
+The general algorithm is the most versatile one as it works with any joint distribution of $\theta$ as long as one can obtain random draws from it.
+As such, this algorithm can be useful when $P(\theta|\mathbf{d})$ does not have a tractable form.
+However, the finite size of the sample adds another layer of uncertainty, potentially requiring significantly more computation time to achieve the same numerical accuracy of the other two algorithms.
 How can EVSI analysis inform study design in predictive analytics?
@@
 In decision-analytic (health policy) modeling, EVSI is typically in net monetary units, and when scaled to population, can be compared with the budget of a planned data collection activity.
-The optimal sample size will be one that maximizes the difference between population EVSI and study costs29.
+The optimal sample size will be one that maximizes the difference between population EVSI and study costs36.
 The NB for risk prediction models, on the other hand, is in net true or false positive units, and as such cannot directly be compared with the budget of a validation study.
@@
 We also do not claim the algorithms we proposed are the only ones that can be used for EVSI computations.
-Other algorithms, such as those based on central limit theorem31,32, can prove useful.
-VoI analysis in decision modeling has received significant boost in computational speed in recent years due to the arrival of algorithms based on non-parametric regression modeling33,34.
+Other algorithms, such as those based on central limit theorem38,39, can prove useful.
+VoI analysis in decision modeling has received significant boost in computational speed in recent years due to the arrival of algorithms based on non-parametric regression modeling40,41.
 This approach can facilitate VoI analysis in risk prediction as well.
 The EVSI defined in this work is for a single, homogeneous target population, and does not consider heterogeneous settings.
-Consider the hierarchical Bayesian tri-variate model by Wynants et al for NB estimation in multi-center validation studies19 and imagine a new center is interested in validating this model in their local setting.
-The posterior predictive distribution from the hierarchical model for a new center (potentially combined with any local information on the performance of the model) can act as ‘current information’, enabling EVSI calculations.
-This will be valid and informative insofar as the expected NB gain in that center is concerned.
-However, in a hierarchical Bayesian framework, information learned from one cluster propagates to others, and thus conducting a validation study in a new center results in change in the distribution of NB in other centers.
-Such NB gain from learning across clusters in hierarchical settings is not reflected in the EVSI computed for this particular setting.
 VoI metrics and corresponding computation algorithms for multi-center studies should be developed separately.
-Further, during external validation, often a secondary aim is to update the model if its performance turns out to be sub-optimal35.
-Such model revision can take different levels of complexity (from intercept correction to re-estimating regression coefficients36) and one might be interested in the expected yield of a given sample size for such model revision.
+Further, during external validation, often a secondary aim is to update the model if its performance turns out to be sub-optimal42.
+Such model revision can take different levels of complexity43 and one might be interested in the expected yield of a given sample size for such model revision.
 This will get connected to the VoI concepts for model development.
 As stated earlier, how VoI metrics for prediction models should inform objective functions for determining the optimal sample size should be debated by the community in the hope of generating consensus and best practice standards.
-There is an ongoing debate on the appropriateness of conventional, frequentist metrics of uncertainty when reporting NB of a clinical prediction model3–5.
-The same concerns can logically be extended to frequentist method for sample size and power calculations around NB37–39.
+There is an ongoing debate on the appropriateness of conventional metrics of uncertainty when reporting NB of a clinical prediction model7–9.
+The same concerns can logically be extended to frequentist method for sample size and power calculations around NB5,6.
 VoI methodology provides a rigorous, utilitarian response to such controversies.